Mr javier's two students loans total $9000.one loan is at5% simple interest and the other is 6 % simple interest. At the end of the year, Mr Javier owes $492 in interest,what is the amount of each loan?
Answer by stanbon(75887)

Mr javier's two students loans total $9000. one loan is at 5% simple interest and the other is 6 % simple interest. At the end of the year, Mr Javier owes $492 in interest,what is the amount of each
--------------------------------
Equations:
f + s = 9000 dollars
---
0.05f + 0.06s = -492 dollars
---------------------------------
Multiply thru 1st by 5.
Multiply thru 2nd by 100.
-------
5f + 5s = 5*9000
5f + 6s = 49200
----------
Subtract and solve for "s":
s = 4200 (amt. invested at 6%)
---
Solve for "f":
f + s = 9000
f = 4800 (amt. invested at 5%)
=================================
Cheers,
Stan H.
===========